Muffy's Muffins had net income of $2,605. The firm retains 60 percent of net income. During the year, the company sold $535 in common stock. What was the cash flow to shareholders?
Answer #1

Hi,

Cash flow to stockholder = dividend paid - new equity raised

here net income = $2,605

retained income = 60%

so dividend percentage =40%

total dividend = 0.4*2605 = $1042

new equity raised= $535

so total cash flow to shareholders= 1042-535 = $507

Thanks